OpenStudy (anonymous):

What is 3 root of 135? A. 5 3 root of 2 B. 2 3 root of 5 C. 5 3 root of 3 D. 3 3 root of 5 If you need me to translate the answer choices just tell me and if you give me a minute i will tell you what i think

OpenStudy (anonymous):

I think D

OpenStudy (anonymous):

@TheAlgebraHelper

OpenStudy (anonymous):

@Ziferot

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

@butterflydreamer

OpenStudy (anonymous):

@dan815

OpenStudy (thomas5267):

\[ \sqrt[3]{135}=?\\ a)\: 5\sqrt[3]{2}\\ b)\: 2\sqrt[3]{5}\\ c)\: 5\sqrt[3]{3}\\ d)\: 3\sqrt[3]{5} \]

OpenStudy (anonymous):

i said I think D

OpenStudy (thomas5267):

Yes that is correct.

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

sorry if that sounded rude or mean

OpenStudy (thomas5267):

I was just typesetting the thing so that others can help.

OpenStudy (anonymous):

ok thank you can you help with one more?

OpenStudy (thomas5267):

Yes?

OpenStudy (anonymous):

ok i tagged you